OpenStudy (anonymous):

Medal 4 fraternal twins! There are four babies in the womb, the technician is unable to tell the gender of the babies. (a) Find the probability that the quadruplets will all be girls. Express your answer as a fraction or an exact decimal - do not round. (b) What is the probability that the couple will have 3 boys and 1 girl? (c) What is the probability that the youngest child will be male?

OpenStudy (caominhim):

1 = boy 0 = girl possibilities are 0000 0001 0010 0011 0100 0101 0110 0111 1000 1001 1010 1011 1100 1101 1110 1111

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

32 possibilities it will be a girl

OpenStudy (caominhim):

there are 16 total possibilities

OpenStudy (caominhim):

you can also find that number because there are 4 babies with 2 total genders and 4^2 = 16

OpenStudy (anonymous):

ok 16/32 right

OpenStudy (caominhim):

where are you getting 32? 16*2? 2^5?

OpenStudy (anonymous):

yes 16*2 = 32

OpenStudy (anonymous):

it says express as a fraction so this is y i got 32

OpenStudy (caominhim):

don't multiply by 2 because there are 16 possibilities your denominator will be 16 you need to find the numeratior \[\frac{?}{16}\]

OpenStudy (anonymous):

4/16

OpenStudy (caominhim):

4/16 what? which question are you answering with that?

OpenStudy (anonymous):

a)

OpenStudy (anonymous):

no sorry 2/16 because it could be a male or female

OpenStudy (caominhim):

look at the possibilities list, how many of those options contain all girls?

OpenStudy (anonymous):

15

OpenStudy (caominhim):

not quite, if the quadruplets were to be all girls then of the 4 digits, every one would be a 0

OpenStudy (anonymous):

ohh i c

OpenStudy (caominhim):

okay so what's your answer for A?

OpenStudy (anonymous):

1/16

OpenStudy (caominhim):

correct

OpenStudy (caominhim):

now for B, 3 boys and 1 girl, you need three 1's and one 0

OpenStudy (anonymous):

3/16

OpenStudy (caominhim):

close, you missed one

OpenStudy (anonymous):

4/16

OpenStudy (caominhim):

correct. And now for number C we're going to have to do use some inference. This question is not worded well but i think it's asking for how many possibilities will the last child out be a boy?

OpenStudy (anonymous):

ya youngest child to be a boy from the fraternal twins

OpenStudy (caominhim):

you can do this one without counting the possibilities, what is the probability that a child will be a boy?

OpenStudy (anonymous):

1/16

OpenStudy (caominhim):

\[\frac{Gender}{Number Of Genders}\]

OpenStudy (anonymous):

1/2

OpenStudy (caominhim):

correct, so 50% of your possibilities will end with a boy so what fraction will give you 50% if your denominator is 16?

OpenStudy (anonymous):

8/16

OpenStudy (caominhim):

that's your answer, though your teacher will probably appreciate it more if you simplify it

OpenStudy (anonymous):

do all the fractions have to be simplfied?

OpenStudy (caominhim):

generally yes

OpenStudy (anonymous):

so its a) 1/16 b) 1/4 c) 1/8 i simlified all of them

OpenStudy (caominhim):

A and B are correct

OpenStudy (caominhim):

for C your unsimplified number is 8/16

OpenStudy (anonymous):

1/2

OpenStudy (caominhim):

correct, you got them all

OpenStudy (anonymous):

thank you :)